2. AN OBJECT IS THROWN WITH A SPEED OF 60M/SEC IN A DIRECTION 60 O ABOVE THE HORIZONTAL. FIND;

A. THE HEIGHT IT RISES

B. THE TOTAL TIME OF FLIGHT

C. THE HORIZONTAL RANGE.


1
Expert's answer
2021-06-01T10:50:09-0400

a)


"h=\frac{60^2\sin^2{60}}{2(9.8)}=138\ m"

b)


"t=2\frac{60\sin{60}}{(9.8)}=10.6\ s"

c)


"R=\frac{60^2\sin{120}}{(9.8)}=318\ m"